OpenStudy (haleyelizabeth2017):

Determine the polar form of the complex number 2-2i. Express the angle theta in degrees, where 0 <= theta <= 360, and round numerical entries in the answer to two decimal places.

OpenStudy (anonymous):

you need two numbers to write \[2-2i\] as \[r\left(\cos(\theta)+i\sin(\theta)\right)\] namely \(r\) and \(\theta\)

OpenStudy (anonymous):

\(r\) is the easiest, it is \[r=\sqrt{a^2+b^2}\]

OpenStudy (haleyelizabeth2017):

\[r=\sqrt{(2)^2+(-2i)^2}\]correct?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

in \(a+bi\) both \(a\) and \(b\) are real numbers

OpenStudy (haleyelizabeth2017):

oh..so it's just \[\sqrt{2^2+(-2)^2}\]?

OpenStudy (anonymous):

so it is just \[\sqrt{2^2+2^2}\] no \(i\)'s involved it is the distance the number is from \((0,0)\) in the complex plane yeah what you wrote

OpenStudy (haleyelizabeth2017):

okay..

OpenStudy (haleyelizabeth2017):

so sqrt8?

OpenStudy (anonymous):

as for \(\theta\) that is pretty easy in this case if you think about where the number is

OpenStudy (anonymous):

yeah \(\sqrt8\) or maybe \(2\sqrt2\)

OpenStudy (anonymous):



OpenStudy (haleyelizabeth2017):

It will be in decimals so 2.83 XD

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

leave it as \[2\sqrt2\]

OpenStudy (haleyelizabeth2017):

Okay...

OpenStudy (anonymous):

can you tell what \(\theta\) is from the picture?

OpenStudy (haleyelizabeth2017):

Not really....perhaps 315 degrees? >.<

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

both sides are the same, so it is like that 45 - 45 - 90 right triangle, so you got it, \(315^\circ\) if you are working in degrees

OpenStudy (haleyelizabeth2017):

awesome

OpenStudy (anonymous):

you could also say \(-45^\circ\)

OpenStudy (haleyelizabeth2017):

But, the problem has a state it in it's positive form XD

OpenStudy (haleyelizabeth2017):

us*

OpenStudy (anonymous):

make sure this is clear, it is an equality between numbers \[2-2i=2\sqrt2\left(\cos(315)+i\sin(315)\right)\]

OpenStudy (anonymous):

the number on the left is equal to the number on the right if you evaluate the trig functions and multiply both by \(2\sqrt2\) you will get \(2-2i\) for sure

OpenStudy (haleyelizabeth2017):

oh